# toltxlabel

### v2

@@ 2 Defining LMTP effects under competing risks
Assume that at each time point $t=1,\ldots\tau$, we measure $X_{t}=(D_{t},Y_{t},L_{t},A_{t},C_{t})$ on each of $i=1,\ldots,n$ study participants, where $L_{t}$ denotes time-varying covariates, $A_{t}$ denotes a vector of treatments at time $t$, $C_{t}$ denotes an indicator of loss-to-follow-up (equal to one if the unit remains on study at time $t+1$ and equal to zero otherwise), $D_{t}$ denotes an indicator of the competing event occurring on or before time $t$, and $Y_{t}$ denotes an indicator of the event of interest not having occurred on or by time $t$. Baseline covariates are included in $L_{1}$. At the end of the study, we measure the outcome of interest $Y_{\tau+1}$. We let $D_{1}=0$ and $Y_{1}=1$, meaning that participants have not experienced the event of interest nor the competing event at the beginning of the study. We let $R_{t}=\mathds{1}\{D_{t}=0,Y_{t}=1\}$ denote an indicator that neither the event of interest nor the competing event have occurred by time $t$. Assume monotone loss-to-follow-up so that $C_{t}=0$ implies $C_{k}=0$ for all $k>t$, in which case all data become degenerate for $k>t$. In this work, we are interested in a scenario where occurrence of the competing event precludes occurrence of the event of interest. For example, in a study looking to assess the effect of hypertension medications on the likelihood of a stroke, death for reasons other than a stroke is a competing event. In this notation, we have $Y_{t}=0$ if $Y_{t-1}=0$, $D_{t}=1$ if $D_{t-1}=1$, and $Y_{t}=0$ if $D_{t}=1$, since the competing event precludes occurrence of the event of interest. Let $X_{i}=(X_{1,i},\ldots,X_{\tau,i})$ denote the observed data for subject $i$ across all time points $t=1,\ldots\tau$. We let $\mathsf{P}f=\int f(x)\,\mathrm{d}\mathsf{P}(x)$ for a given function $f(x)$. We use $\mathsf{P}_{n}$ to denote the empirical distribution of $X_{1},\ldots\,X_{n}$, and assume $\mathsf{P}$ is an element of the nonparametric statistical model defined as all continuous densities on $X$ with respect to a dominating measure $\nu$. We let $\mathsf{E}$ denote the expectation with respect to $\mathsf{P}$, i.e., $\mathsf{E}\{f(X)\}=\int f(x)\,\mathrm{d}\mathsf{P}(x)$. We also let $||f||^{2}$ denote the $L_{2}(\mathsf{P})$ norm $\int f^{2}(x)\,\mathrm{d}\mathsf{P}(x)$. We use $\bar{W}_{t}=(W_{1},\ldots,W_{t})$ to denote the past history of a variable $W$, use $\underaccent{\bar}{W}_{t}=(W_{t},\ldots,W_{\tau})$ to denote the future of a variable, and use $H_{t}=(\bar{D}_{t},\bar{Y}_{t},\bar{L}_{t},\bar{A}_{t-1})$ to denote the history of the time-varying covariates and the treatment process until just before time $t$. $\bar{W}$ denotes the entire history $(W_{1},\ldots,W_{\tau})$ of a variable. By convention, variables with an index $t\leq 0$ are defined as the null set, expectations conditioning on a null set are marginal, products of the type $\prod_{t=k}^{k-1}b_{t}$ and $\prod_{t=0}^{0}b_{t}$ are equal to one, and sums of the type $\sum_{t=k}^{k-1}b_{t}$ and $\sum_{t=0}^{0}b_{t}$ are equal to zero. For vectors $u$ and $v$, $u\leq v$ denotes point-wise inequalities.
@@
[…]
@@
[…]
@@ 2.1 Identification of the effect of LMTPs under competing risks
Set $\mathsf{q}_{\tau+1}=Y_{\tau+1}$ and $R_{\tau+1}=1$. For $t=\tau,\ldots,1$, recursively define
$\mathsf{q}_{t}:(a_{t},h_{t})\mapsto\mathsf{E}\left[R_{t+1}\times\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\right],$ (5)
@@ 3 Efficient estimation via sequentially doubly robust regression
$\varphi_{t}:x\mapsto\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{w}_{k}(c_{k},a_{k},h_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{d}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$ (8)
@@
$\mathsf{E}\big{[}R_{t+1}\times\varphi_{t+1}(X;\eta^{\prime})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}=\mathsf{q}_{t}(a_{t},h_{t}).$
@@
Consider the odds ratio incremental propensity score interventions of Example 3, where we let $\varepsilon_{t}\sim U(0,1)$, and define $\mathbbm{d}(h_{t},\varepsilon_{t})=\mathds{1}\{\varepsilon_{t}<\mathsf{g}_{t}^{\mathbbm{d}}(1\mid h_{t})\}$. The result in Proposition 2 also holds for this intervention, but it does not allow for doubly robust estimation. To see why, notice that $\mathsf{q}_{t}$ itelse depends on $\{\mathsf{g}_{s}:s>t\}$ (or equivalently, on $\{\mathsf{w}_{s}:s>t\}$) through $\mathbbm{d}$. Therefore, if $\mathsf{q}^{\prime}_{s}=\mathsf{q}_{s}$ and $\mathsf{w}^{\prime}_{s}\neq\mathsf{w}_{s}$ for all $s>t$, we get $\mathsf{E}\big{[}R_{t+1}\times\varphi_{t+1}(X;\eta^{\prime})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}=\mathsf{q}_{t}(a_{t},h_{t})$, where $\mathsf{q}_{t}$ is evaluated at the misspecified $\{\mathsf{w}_{s}^{\prime}:s>t\}$. This is aligned with the findings of Kennedy (2019). In contrast, the risk ratio incremental propensity score intervention does allow for doubly robust estimation, since $\mathbbm{d}_{t}$ does not depend on $\mathsf{w}_{t}$.
@@ 2. Step 2:
    • Compute the pseudo-outcome $\check{Y}_{t+1,i}=R_{t+1}\times\varphi_{t+1}(X_{i};\underline{\check{\eta}}_{t,j(i)})$ for all $i=1,\ldots,n$.
@@ 3. Step 3:
$\check{\mathsf{q}}_{t}^{\dagger}(a_{t},h_{t})=\mathsf{E}\big{[}R_{t+1}\varphi_{t+1}(X;\underline{\check{\eta}}_{t})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]},$
@@
$\mathsf{q}_{t}^{\star}(a_{t},h_{t})=\mathsf{E}\big{[}R_{t+1}\hat{\mathsf{q}}_{t+1}(a_{t+1},h_{t+1})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}.$
@@ Appendix B Identification (Proposition 1)
$=\mathds{1}\{d_{t}=0,y_{t}=1\}\mathsf{E}[Y_{\tau+1}\mid D_{t}=0,Y_{t}=C_{t}=1,H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}}]$
$=r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t}).$
@@
$=\mathsf{E}[R_{t+1}\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})\mid D_{t}=d_{t},Y_{t}=y_{t},H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}},C_{t}=1]$
$=r_{t}\mathsf{E}[R_{t+1}\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})\mid R_{t}=C_{t}=1,H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}}]$
$=r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t}).$
Since $R_{1}=1$, we have $\mathsf{m}_{1}(E_{1}^{\mathbbm{h}},V_{1})=\mathsf{q}_{1}(A_{1}^{\mathbbm{d}},H_{1})$, concluding the proof of the result. ∎
@@ Appendix C Efficient influence function
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-r_{s}\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)r_{s}\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})r_{k}\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$
$=r_{t}\left[\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{w}_{t}(c_{k},a_{k},h_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})\right]$
$=r_{t}\varphi_{t}(x).$ (10)
An application of Theorem 3 together with the above results for $t=1$ yields the desired result after noticing that $R_{1}=1$. ∎
@@
$\mathsf{q}_{t}(a_{t},h_{t})=\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}]+\mathsf{J}_{t}(a_{t},h_{t}).$
@@
$=\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)\mid D_{t}=d_{t},Y_{t}=y_{t},C_{t}=c_{t},A_{t}=a_{t},H_{t}=h_{t}]+\mathsf{U}_{t}(e_{t},v_{t})$
$=r_{t}\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}]+r_{t}\mathsf{J}_{t}(a_{t},h_{t}),$

### v3

@@ 2 Defining LMTP effects under competing risks
Assume that at each time point $t=1,\ldots\tau$, we measure $X_{t}=(D_{t},Y_{t},L_{t},A_{t},C_{t})$ on each of $i=1,\ldots,n$ study participants, where $L_{t}$ denotes time-varying covariates, $A_{t}$ denotes a vector of treatments at time $t$, $C_{t}$ denotes an indicator of loss-to-follow-up (equal to one if the unit remains on study at time $t+1$ and equal to zero otherwise), $D_{t}$ denotes an indicator of the competing event occurring on or before time $t$, and $Y_{t}$ denotes an indicator of the event of interest not having occurred on or by time $t$. Baseline covariates are included in $L_{1}$. At the end of the study, we measure the outcome of interest $Y_{\tau+1}$. We let $D_{1}=0$ and $Y_{1}=1$, meaning that participants have not experienced the event of interest nor the competing event at the beginning of the study. We let $R_{t}=\mathds{1}\{D_{t}=0,Y_{t}=1\}$ denote an indicator that neither the event of interest nor the competing event have occurred by time $t$, and $Z_{t}=\mathds{1}\{D_{t}=1,Y_{t}=1\}$ denote an indicator that the event of interest has not occurred but the competing event has. Assume monotone loss-to-follow-up so that $C_{t}=0$ implies $C_{k}=0$ for all $k>t$, in which case all data become degenerate for $k>t$. In this work, we are interested in a scenario where occurrence of the competing event precludes occurrence of the event of interest. For example, in a study looking to assess the effect of hypertension medications on the likelihood of a stroke, death for reasons other than a stroke is a competing event. In this notation, we have $Y_{t}=0$ if $Y_{t-1}=0$, $D_{t}=1$ if $D_{t-1}=1$, and $Y_{t}=1$ if $D_{t}=1$ and $Y_{t-1}=1$, since the competing event precludes occurrence of the event of interest. Let $X_{i}=(X_{1,i},\ldots,X_{\tau,i})$ denote the observed data for subject $i$ across all time points $t=1,\ldots\tau$. We let $\mathsf{P}f=\int f(x)\,\mathrm{d}\mathsf{P}(x)$ for a given function $f(x)$. We use $\mathsf{P}_{n}$ to denote the empirical distribution of $X_{1},\ldots\,X_{n}$, and assume $\mathsf{P}$ is an element of the nonparametric statistical model defined as all continuous densities on $X$ with respect to a dominating measure $\nu$. We let $\mathsf{E}$ denote the expectation with respect to $\mathsf{P}$, i.e., $\mathsf{E}\{f(X)\}=\int f(x)\,\mathrm{d}\mathsf{P}(x)$. We also let $||f||^{2}$ denote the $L_{2}(\mathsf{P})$ norm $\int f^{2}(x)\,\mathrm{d}\mathsf{P}(x)$. We use $\bar{W}_{t}=(W_{1},\ldots,W_{t})$ to denote the past history of a variable $W$, use $\underaccent{\bar}{W}_{t}=(W_{t},\ldots,W_{\tau})$ to denote the future of a variable, and use $H_{t}=(\bar{D}_{t},\bar{Y}_{t},\bar{L}_{t},\bar{A}_{t-1})$ to denote the history of the time-varying covariates and the treatment process until just before time $t$. $\bar{W}$ denotes the entire history $(W_{1},\ldots,W_{\tau})$ of a variable. By convention, variables with an index $t\leq 0$ are defined as the null set, expectations conditioning on a null set are marginal, products of the type $\prod_{t=k}^{k-1}b_{t}$ and $\prod_{t=0}^{0}b_{t}$ are equal to one, and sums of the type $\sum_{t=k}^{k-1}b_{t}$ and $\sum_{t=0}^{0}b_{t}$ are equal to zero. For vectors $u$ and $v$, $u\leq v$ denotes point-wise inequalities.
@@
$\mathbbm{d}(a_{t},\varepsilon_{t})=\begin{cases}a_{t}&\text{if }\varepsilon_{t}<\delta\\ 0&\text{otherwise},\end{cases}$ (3)
@@
$\mathbbm{d}(a_{t},h_{t})=\begin{cases}1&\text{if }l^{\prime}_{t-m}=1\\ a_{t}&\text{otherwise}.\end{cases}$ (4)
@@ 2.1 Identification of the effect of LMTPs under competing risks
Set $\mathsf{q}_{\tau+1}=Y_{\tau+1}$, $R_{\tau+1}=1$ and $Z_{\tau+1}=0$. For $t=\tau,\ldots,1$, recursively define
$\mathsf{q}_{t}:(a_{t},h_{t})\mapsto\mathsf{E}\left[R_{t+1}\times\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})+Z_{t+1}\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\right],$ (5)
@@ 3 Efficient estimation via sequentially doubly robust regression
$\varphi_{t}:x\mapsto\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{w}_{k}(c_{k},a_{k},h_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{d}},h_{s+1})+z_{s+1}-\mathsf{q}_{s}(a_{s},h_{s})\}+\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$ (8)
@@
$\mathsf{E}\big{[}R_{t+1}\times\varphi_{t+1}(X;\eta^{\prime})+Z_{t+1}\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}=\mathsf{q}_{t}(a_{t},h_{t}).$
@@
Consider the odds ratio incremental propensity score interventions of Example 3, where we let $\varepsilon_{t}\sim U(0,1)$, and define $\mathbbm{d}(h_{t},\varepsilon_{t})=\mathds{1}\{\varepsilon_{t}<\mathsf{g}_{t}^{\mathbbm{d}}(1\mid h_{t})\}$. The result in Proposition 2 also holds for this intervention, but it does not allow for doubly robust estimation. To see why, notice that $\mathsf{q}_{t}$ itelse depends on $\{\mathsf{g}_{s}:s>t\}$ (or equivalently, on $\{\mathsf{w}_{s}:s>t\}$) through $\mathbbm{d}$. Therefore, if $\mathsf{q}^{\prime}_{s}=\mathsf{q}_{s}$ and $\mathsf{w}^{\prime}_{s}\neq\mathsf{w}_{s}$ for all $s>t$, we get $\mathsf{E}\big{[}R_{t+1}\times\varphi_{t+1}(X;\eta^{\prime})+Z_{t+1}\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}=\mathsf{q}_{t}(a_{t},h_{t})$, where $\mathsf{q}_{t}$ is evaluated at the misspecified $\{\mathsf{w}_{s}^{\prime}:s>t\}$. This is aligned with the findings of Kennedy (2019). In contrast, the risk ratio incremental propensity score intervention does allow for doubly robust estimation, since $\mathbbm{d}_{t}$ does not depend on $\mathsf{w}_{t}$.
@@ 2. Step 2:
    • Compute the pseudo-outcome $\check{Y}_{t+1,i}=R_{t+1}\times\varphi_{t+1}(X_{i};\underline{\check{\eta}}_{t,j(i)})+Z_{t+1}$ for all $i=1,\ldots,n$.
@@ 3. Step 3:
$\check{\mathsf{q}}_{t}^{\dagger}(a_{t},h_{t})=\mathsf{E}\big{[}R_{t+1}\varphi_{t+1}(X;\underline{\check{\eta}}_{t})+Z_{t+1}\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]},$
@@
$\mathsf{q}_{t}^{\star}(a_{t},h_{t})=\mathsf{E}\big{[}R_{t+1}\hat{\mathsf{q}}_{t+1}(a_{t+1},h_{t+1})+Z_{t+1}\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}.$
@@ Appendix B Identification (Proposition 1)
$=\mathds{1}\{d_{t}=0,y_{t}=1\}\mathsf{E}[Y_{\tau+1}\mid D_{t}=0,Y_{t}=C_{t}=1,H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}}]+\mathds{1}\{d_{t}=1,y_{t}=1\}$
$=r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})+z_{t}.$
@@
$=\mathsf{E}[R_{t+1}\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})+Z_{t+1}\mid D_{t}=d_{t},Y_{t}=y_{t},H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}},C_{t}=1]$
$=r_{t}\mathsf{E}[R_{t+1}\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})+Z_{t+1}\mid R_{t}=C_{t}=1,H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}}]+z_{t}$
$=r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})+z_{t}.$
Since $R_{1}=1$ and $Z_{1}=0$, we have $\mathsf{m}_{1}(E_{1}^{\mathbbm{h}},V_{1})=\mathsf{q}_{1}(A_{1}^{\mathbbm{d}},H_{1})$, concluding the proof of the result. ∎
@@ Appendix C Efficient influence function
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})+z_{s+1}-r_{s}\mathsf{q}_{s}(a_{s},h_{s})-z_{s}\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})+z_{t}$
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)r_{s}\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})+z_{s+1}-\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})+z_{t}$
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})r_{k}\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})+z_{s+1}-\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})+z_{t}$
$=r_{t}\left[\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{w}_{t}(c_{k},a_{k},h_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})+z_{s+1}-\mathsf{q}_{s}(a_{s},h_{s})\}+\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})\right]+z_{t}$
$=r_{t}\varphi_{t}(x)+z_{t},$ (10)
where we used that $r_{s}=0$ implies $r_{s+1}=0$ and $z_{s}=z_{s+1}$, and $r_{s}=1$ implies $z_{s}=0$. An application of Theorem 3 together with the above results for $t=1$ yields the desired result after noticing that $R_{1}=1$. ∎
@@
$\mathsf{q}_{t}(a_{t},h_{t})=\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)+Z_{t+1}\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}]+\mathsf{J}_{t}(a_{t},h_{t}).$
@@
$=\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)+Z_{t+1}\mid D_{t}=d_{t},Y_{t}=y_{t},C_{t}=c_{t},A_{t}=a_{t},H_{t}=h_{t}]+\mathsf{U}_{t}(e_{t},v_{t})$
$=r_{t}\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)+Z_{t+1}\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}]+z_{t}+r_{t}\mathsf{J}_{t}(a_{t},h_{t}),$